# Prompt-based Distribution Alignment for Unsupervised Domain Adaptation

### 15 Dec 2023 v1

@@ 3.2 Revisiting Prompt Learning
$p(y=i\mid x)=\frac{\exp\left(sim\left(w_{i},z)\right)/\tau\right)}{\sum^{K}_{j=1}\exp\left(sim\left(w_{j},z)\right)/\tau\right)},$ (1)
where $\tau$ denotes temperature parameter, $K$ denotes the number of classes and $sim$ denotes the cosine similarity.
@@ 4.1 Prompting for Base Branch
$\mathcal{L}_{u}=-\mathbb{I}{(\tau_{p}\geq\tau)}\sum_{i}\hat{y}^{t}_{i}log\frac{\exp(sim(\hat{w}_{i},\hat{z}^{t})/\tau)}{\sum^{K}_{j=1}\exp(sim(\hat{w}_{j},\hat{z}^{t})/\tau)},$ (3)
@@ 4.2 Pipeline of Alignment Branch
$\mathcal{L}_{ua}=-\mathbb{I}{(\tau_{p}\geq\tau)}\sum_{i}\hat{y}^{t}_{i}log\frac{\exp(sim(\hat{w}_{i},h(\hat{z}^{t}))/\tau)}{\sum^{K}_{j=1}\exp(sim(\hat{w}_{j},h(\hat{z}^{t}))/\tau)}.$ (8)

### 26 Jan 2024 v2

@@ 3.2 Revisiting Prompt Learning
$p(y=i\mid x)=\frac{\exp\left(sim\left(w_{i},z)\right)/t\right)}{\sum^{K}_{j=1}\exp\left(sim\left(w_{j},z)\right)/t\right)},$ (1)
where $t$ denotes temperature parameter, $K$ denotes the number of classes and $sim$ denotes the cosine similarity.
@@ 4.1 Prompting for Base Branch
$\mathcal{L}_{u}=-\mathbb{I}{(\tau_{p}\geq\tau)}\sum_{i}\hat{y}^{t}_{i}log\frac{\exp(sim(\hat{w}_{i},\hat{z}^{t})/t)}{\sum^{K}_{j=1}\exp(sim(\hat{w}_{j},\hat{z}^{t})/t)},$ (3)
@@ 4.2 Pipeline of Alignment Branch
$\mathcal{L}_{ua}=-\mathbb{I}{(\tau_{p}\geq\tau)}\sum_{i}\hat{y}^{t}_{i}log\frac{\exp(sim(\hat{w}_{i},h(\hat{z}^{t}))/t)}{\sum^{K}_{j=1}\exp(sim(\hat{w}_{j},h(\hat{z}^{t}))/t)}.$ (8)